the income of A and B are in the ratio of 7:5. their expenses are In the ratio of 9:8.an increase of A’s expenditure by rupees 100 and decrease of B’s expenditure by rupees 200 makes their savings rupees 400 each. Find their incomes
Answer by dkppathak(439)

the income of A and B are in the ratio of 7:5. their expenses are In the ratio of 9:8.an increase of A’s expenditure by rupees 100 and decrease of B’s expenditure by rupees 200 makes their savings rupees 400 each. Find their incomes
solution
let income is 7x and 5x respectively for A and B
expenditure for A and B is 9Y and 8Y
as per given conditions
new a expenditure is 9Y+100 and B is 8Y-200
saving for both is now 400
A saving=7x-(9Y+100)=400
B saving =5X-(8Y-200)=400
7x-9y-100=400 or 7x-9y=500 (1)
5x-8y+200=400 or 5x-8y=200 (2) by solving these equations by elimination methods Y=100 and x= 200
income of A=1400 and B=1000